OpenStudy (anonymous):

use the substitution method to find the value of y in the given system of equations.express non-integer values as fractions in lowest terms. y=2x+5 x+y=4

OpenStudy (anonymous):

Put y = 2x + 5 in second equation : x + 2x + 5 = 4 3x = -1 x = -1/3 and then put this value in x + y = 4 to get the value of y.. Getting or not??

OpenStudy (anonymous):

yupp i get it thank you! i got 13/3 for the vaule of y is that correct?